1/2

John Ellerton,

The day Thou gavest, Lord, is ended

1. The day Thou gavest, Lord, is ended, The darkness falls at Thy behest; To Thee our morning hymns ascended, Thy praise shall sanctify our rest.

2. We thank Thee that Thy church, unsleeping, While earth rolls onward into light, Through all the world her watch is keeping, And rests not now by day or night.

3. As o’er each continent and island The dawn leads on another day, The voice of prayer is never silent, Nor dies the strain of praise away.

4. The sun that bids us rest is waking Our brethren ’neath the western sky, And hour by hour fresh lips are making Thy wondrous doings heard on high.

5. So be it, Lord; Thy throne shall never, Like earth’s proud empires, pass away: Thy kingdom stands, and grows forever, Till all Thy creatures own Thy sway.

…

2/2

Verkinto de tiu ĉi Angla poemo estas John Ellerton (∗1826-12-16 – †1893-06-15), publikigita 1870. Arg-130-255 (2008-11-08 14:47:39) Information about the author can be found on the website http: // www. cyberhymnal. org/ bio/ e/ l/ l/ ellerton_j. htm .

http://www.poezio.net/version?poem-id=130&version-id=,255